# Gravitational waves from thermal heavy scalar dark matter

### v1

@@ Abstract
We propose a gauge singlet scalar with mass around $1-10$ TeV as a thermal heavy dark matter candidate along with a dilaton as a Higgs portal mediator in a dimensionless scalar extension of the Standard Model. We demonstrate analytically that such a model gives rise to very strong first-order electroweak phase tyransition through supercooling. We calculate the corresponding gravitational wave signals due to bubble collisions during the phase transition. The produced gravitational waves can detected by future space-based gravitational wave detectors in the frequency range from $10^{-4}$ Hz to $0.1$ Hz.
@@ 1 Introduction
After the discovery of first spin zero scalar elementary particle at the LHC in 2012 [1, 2, 3], it has been a reasonable assumption to expect for more scalars in nature. While the existence of other types of particles are possible in beyond the standard model (BSM), but an addition of even one scalar degree of freedom apart from the Higgs particle, complicates significantly the process of the electroweak symmetry breaking [4, 5]. The scalar extensions of the standard model have been always of interest to explain the theoretical and observational shortcomings of the standard model (SM). Additional scalars may accommodate the dark matter problem [6, 7, 8, 9, 10], or to impose the electroweak symmetry breaking in the early universe be strongly first order for a successful explanation of the observed matter-anti matter asymmetry in the universe [11], through which gravitational wave signals can be detected [12, 13], or to address a combination of the aforementioned problems [14, 15, 16].
Additional scalars can be employed in dimentionless extensions of the SM to address additionally the heirarchy and fine-tuning problems. There are scale invariant models that have tried to explain one or more problems such as dark matter or electroweak phase transition [17, 18, 19, 20, 21, 22]. Although it is known that electroweak phase transition in scale invariant scalar extension of the SM is very strong [19], but the possibility of detecting the gravitational waves signals produced during such phase transitions when an extra scalar plays the role of a thermal heavy dark matter has not been investigated in the literature. A nonthermal heavy dark matter scenario from first-order phase transition is studied in [23].
@@ 2 Model
The scale symmetry breaking which subsequently leads to electroweak symmetry breaking takes place through radiative corrections. According to Gildener-Weinberg approach there exists a flat direction in the field space along which the potential and its minimum is vanishing. In our model, the flat direction with three scalar field is three-dimentional $(n_{1},n_{2},n_{3})$. As the third scalar, i.e. $\phi_{3}$ is the DM candidate here, it does not acquire a non-zero VEV, therefore we set $n_{3}=0$ and the flat direction will be in $(\phi_{1},\phi_{2})$ space.
@@
$B=\frac{1}{64\pi^{2}\braket{\phi}^{4}}\left(+m_{DM}^{4}m_{H}^{4}+6m_{W}^{4}+3m_{Z}^{4}-12m_{t}^{4}\right)$ (6)
@@ 4 Electroweak Phase Transition
We assume that the scale symmetry is broken at high enough temperatures respect to the masses in the model so that we can use the high-temperature expansion. The thermal effective potential in the high-temperature approximation is given by [35]
@@ 5 Gravitational Waves
$36.93$ TeV $2.8$ TeV $117$ GeV $1.1\times 10^{5}$ $252$ $72.27\times 10^{-}$ 7 $11.32\times 10^{-}$ 1 $1.22\times 10^{-1}$
116.64×116.64$\times$116.64 × TeV $8.86$ TeV $177$ GeV $2.14\times 10^{6}$ GeV $4.52$ $-7\times 10^{-9}$ $4521$ $1.36\times 10^{-1}$
81.85×81.85$\times$81.85 × TeV $6.2$ TeV $47$ GeV $2.14\times 10^{6}$ GeV $1190$ $-7\times 10^{-9}$ $4.53$ $1.36\times 10^{-1}$
Table 1: Shown are benchmarks for strongly first-order electroweak phase transition and gravitational waves parameters while the DM relic density is chosen to be the observed value.
Gravitational waves might stem from strong first-order phase transition in early universe in different ways (see e.g. [37]): 1) the bubble walls collisions and schocks in the plasma, $\Omega_{\phi}h^{2}$, which is the contribution of the scalar field $\phi$ using a technique known as ’envelope approximation’, 2) contributions from the sound waves, $\Omega_{\text{sw}}h^{2}$, produced by the bubble wall collision and 3) from the magnetohydrodynamic (MHD) turbulence in the plasma $\Omega_{\text{tur}}$. The total stochastic GW background is approximately a linear combination of all contributions,
@@
In table 1, we have chosen from lightest to heaviest DM mass possible with corresponding parameters which give rise to the correct observed DM relic abundance. For the DM mass $m_{\text{DM}}\sim 350$ GeV the coupling $\lambda_{12}$ is of order $210^{-}$ 2 while for the heaviest DM mass in the table, $m_{\text{DM}}\sim 100$ TeV the coupling $\lambda_{12}$ becomes as small as $910^{-}$ 9. For the parameter $\alpha$ that is opposite; the heavier the DM mass is, the larger parameter $\alpha$ becomes, i.e. from $\alpha\sim 3$ to $\alpha\sim 10^{6}$. Another point in table 1 is that for heavy DM there is a suppercooling in electroweak phase transitions, which is of course expected from the large values for the $\alpha$ parameter. Although the critical temperature lies around TeV scale, the nucleation temperature is always around the electroweak scale. For the benchmarks given in table 1 we have plotted the gravitational waves signal produced during the bubble wall collisions in terms of the nucleation temperature $T_{n}$ and the ratio $\beta/H_{*}$. The result is shown in figure 3. We see that the for heavy dark matter candidates in the current model, the GW signals can be detected only by AEDGE+, AEDGE and LISA with GW frequency from around $10^{-4}$ Hz to $0.2$ Hz.

### v3

@@ Abstract
We propose a gauge singlet scalar with mass around $1-100$ TeV as a thermal heavy dark matter candidate along with a dilaton as a Higgs portal mediator in a dimensionless scalar extension of the Standard Model. We demonstrate analytically that such a model gives rise to very strong first-order electroweak phase transition through supercooling. We calculate the corresponding gravitational wave signals due to bubble collisions during the phase transition. The produced gravitational waves can be detected by future space-based gravitational wave detectors in the frequency range from $10^{-4}$ Hz to $0.1$ Hz.
@@ 1 Introduction
After the discovery of first spin zero scalar elementary particle at the LHC in 2012 [1, 2, 3], it has been a reasonable assumption to expect more scalars in nature. While the existence of other types of particles is possible in beyond the standard model (BSM), but an addition of even one scalar degree of freedom apart from the Higgs particle, complicates significantly the process of the electroweak symmetry breaking [4, 5]. The scalar extensions of the standard model have been always of interest to explain the theoretical and observational shortcomings of the standard model (SM). Additional scalars may accommodate the dark matter problem [6, 7, 8, 9, 10], or to impose the electroweak symmetry breaking in the early universe be strongly first order for a successful explanation of the observed matter-anti matter asymmetry in the universe [11], through which gravitational wave signals can be detected [12, 13], or to address a combination of the aforementioned problems [14, 15, 16].
Additional scalars can be employed in dimensionless extensions of the SM to address additionally the hierarchy and fine-tuning problems. There are scale invariant models that have tried to explain one or more problems such as dark matter or electroweak phase transition [17, 18, 19, 20, 21, 22, 23, 24]. Although it is known that electroweak phase transition in scale invariant scalar extension of the SM is very strong [19], but the possibility of detecting the gravitational waves signals produced during such phase transitions when an extra scalar plays the role of a thermal heavy dark matter has not been investigated in the literature. A nonthermal heavy dark matter scenario from first-order phase transition is studied in [25].
@@ 2 Model
The scale symmetry breaking which subsequently leads to electroweak symmetry breaking takes place through radiative corrections. According to Gildener-Weinberg approach there exists a flat direction in the field space along which the potential and its minimum is vanishing. In our model, the flat direction with three scalar field is three-dimensional $(n_{1},n_{2},n_{3})$. As the third scalar, i.e. $\phi_{3}$ is the DM candidate here, it does not acquire a non-zero VEV, therefore we set $n_{3}=0$ and the flat direction will be in $(\phi_{1},\phi_{2})$ space.
@@
$B=\frac{1}{64\pi^{2}\braket{\phi}^{4}}\left(m_{DM}^{4}+m_{H}^{4}+6m_{W}^{4}+3m_{Z}^{4}-12m_{t}^{4}\right)$ (6)
@@ 4 Electroweak Phase Transition
We assume that the scale symmetry is broken at high enough temperatures respect to the masses in the model so that we can use the high-temperature expansion.  [^1] If $m/T\lesssim 1.6$, the high-temperature approximation agrees better than $5\%$ with the exact thermal potential. For large values of $m/T$, still the high-temperature approximation is valid up to $10\%$ due to suppressed Boltzmann contribution (see appendix B in [37]). The thermal effective potential in the high-temperature approximation is given by [38]
@@ 5 Gravitational Waves
$36.93$ TeV $2.8$ TeV $117$ GeV $1.1\times 10^{5}$ $252$ $-2.27\times 10^{-7}$ $1.32\times 10^{-1}$ $1.22\times 10^{-1}$
$81.85$ TeV $6.21$ TeV $47$ GeV $1.08\times 10^{8}$ $1190$ $-3.11\times 10^{-8}$ $1.73$ $2.67\times 10^{-2}$
$116.64$ TeV $8.86$ TeV $177$ GeV $2.14\times 10^{6}$ $4521$ $-7\times 10^{-9}$ $4.53$ $1.36\times 10^{-2}$
Table 1: Shown are benchmarks for strongly first-order electroweak phase transition and gravitational wave parameters. For the four last benchmarks, the observed DM relic density cannot be fully accounted by the model due to DM dilution following the phase transition.
Gravitational waves might stem from strong first-order phase transition in early universe in different ways (see e.g. [40]): 1) the bubble walls collisions and shocks in the plasma, $\Omega_{\phi}h^{2}$, which is the contribution of the scalar field $\phi$ using a technique known as ’envelope approximation’, 2) contributions from the sound waves, $\Omega_{\text{sw}}h^{2}$, produced by the bubble wall collision and 3) from the magnetohydrodynamic (MHD) turbulence in the plasma $\Omega_{\text{tur}}$. The total stochastic GW background is approximately a linear combination of all contributions,
@@
In table 1, we have chosen from lightest to heaviest DM mass possible with corresponding parameters which give rise to the correct observed DM relic abundance. For the DM mass $m_{\text{DM}}\sim 350$ GeV the coupling $\lambda_{12}$ is of order $10^{-2}$ while for the heaviest DM mass in the table, $m_{\text{DM}}\sim 100$ TeV the coupling $\lambda_{12}$ becomes as small as $10^{-9}$. For the parameter $\alpha$ that is opposite; the heavier the DM mass is, the larger parameter $\alpha$ becomes, i.e. from $\alpha\sim 3$ to $\alpha\sim 10^{6}$ For large $\alpha$, due to substantial amount of latent heat and therefore a large entropy injection into the SM sector, the relic density will be diluted [47]. Taking into account this effect, the four last benchmarks in Table 1, cannot account for all observed DM relic abundance. Another point in table 1 is that for heavy DM there is a suppercooling in electroweak phase transitions, which is of course expected from the large values for the $\alpha$ parameter. Although the critical temperature lies around TeV scale, the nucleation temperature is always around the electroweak scale. For the benchmarks given in table 1 we have plotted the gravitational waves signal produced during the bubble wall collisions in terms of the nucleation temperature $T_{n}$ and the ratio $\beta/H_{*}$. The result is shown in figure 3. We see that the for heavy dark matter candidates in the current model, the GW signals can be detected only by AEDGE+, AEDGE and LISA with GW frequency from around $10^{-4}$ Hz to $0.2$ Hz. This is compatible with the results in [48].